Solve each of the following systems by addition. If a unique solution does not exist, state
whether the system is inconsistent or dependent.
x + 5y = 10
-2x - 10y = -20

Answer by jainenderkapoor(61)

x + 5y = 10
-2x - 10y = -20
If I look at the equations carefully...
I see that I can get the (2) equation by simply multiplying (1) by -2
These type of equations are called dependent equations, as they will have same solution. If we try to plot a graph for each of the equation.. we will get only one single line..
Please visit -- http://www.purplemath.com/modules/systlin2.htm -- to see and learn more about system of linear equations.

Hope the solution is clear to you. In case you have any other doubt, you are welcome to contact me on jainenderkapoor@gmail.com I am also available for Online Tutoring at very reasonable rates.
Jainender Kapoor